# Mixture of Experts in Image Classification: What’s the Sweet Spot?

### v1

@@ Abstract
Mixture-of-Experts (MoE) models have shown promising potential for parameter-efficient scaling across various domains. However, the implementation in computer vision remains limited, and often requires large-scale datasets comprising billions of samples. In this study, we investigate the integration of MoE within computer vision models and explore various MoE configurations on open datasets. When introducing MoE layers in image classification, the best results are obtained for models with a moderate number of activated parameters per sample. However, such improvements gradually vanish when the number of parameters per sample increases.
Introduction
@@
Recent advancements in machine learning, particularly in the domains of natural language processing  (NLP) (Vaswani et al. 2017; Kenton and Toutanova 2019) and computer vision (Dosovitskiy et al. 2020), have been primarily driven by scaling up model size, computational budgets, and training data. While these large-scale models demonstrate impressive performance, they are often expensive to train and consume considerable energy resources (Strubell, Ganesh, and McCallum 2019). As a result, the research community has been increasingly interested in exploring more efficient training and serving paradigms. One such promising solution is the use of sparse expert models, with Mixture-of-Experts (MoE) (Shazeer et al. 2017; Lepikhin et al. 2020).
@@
MoE models introduce sparsity by partitioning the set of parameters into multiple parallel sub-models, referred to as “experts”. During training and inference, the gating part of the models routes input examples to specific expert(s), ensuring that each example only interacts with a subset of the network parameters. As the computational cost is, partially, correlated with the number of parameters activated for a given sample rather than the total number of parameters, this approach facilitates the scaling-up of the model, while keeping computational costs under control, making it an attractive option for a wide range of applications.
@@
Despite their success in various domains (Hihn and Braun 2021; Costa-jussà et al. 2022; Zoph et al. 2022; Fedus, Zoph, and Shazeer 2022), the application of MoE models in computer vision remains limited, and it often requires very large datasets (Riquelme et al. 2021; Mustafa et al. 2022; Komatsuzaki et al. 2022) to be competitive against state-of-the-art approaches. In this work, we first focus on leveraging the potential of MoE models for computer vision on ImageNet-1k and ImageNet-21k (Russakovsky et al. 2015). We study the efficiency of integrating MoE within two renowned architectures, ConvNext (Liu et al. 2022) and Vision Transformer (ViT) (Touvron, Cord, and Jégou 2022). We conduct a series of experiments considering various architecture configurations. Likewise, we investigate the impact of various components, including the number of experts and their sizes, the gate design, and the layer positions, among others. Our experimental findings indicate that the optimal design is contingent on the specific network architecture, and that consistently situating the MoE layer within the final two even blocks invariably yields substantial improvements for moderate model size. However when scaling up the approach to large models and large datasets, close to the state of the art, the benefits of using MoE for image classification seem to gradually vanish.
Related Works
@@
The Mixture-of-Experts (MoE) model, introduced by Jacobs et al. (1991), partitions complex tasks into hopefully simpler sub-tasks handled by expert models, whose predictions are combined to produce the final output. This framework has been successfully integrated into various neural network architectures, particularly transformers for NLP tasks (Shazeer et al. 2017). Given this success in NLP, interest has burgeoned in applying MoE to the diverse and complex realm of computer vision.
@@
The transformative potential of MoE architectures is underscored by their successful integration across various domains, with the Vision Transformer (ViT) being a prime example. Riquelme et al. (2021) introduced V-MoE, an MoE-augmented ViT model, for image classification tasks on a massive dataset containing hundreds of millions of examples. They showed that V-MoE does not only match the performance of prior state-of-the-art architectures, but it also requires half the computational resources during inference. Lou et al. (2021) presented a sparse MoE MLP model based on the MLP-Mixer architecture (Tolstikhin et al. 2021). While the MLP-Mixer architecture does not display performance accuracy, the MoE-enhanced version surpassed its dense counterpart in experiments on ImageNet and CIFAR. Hwang et al. (2022) demonstrated the effectiveness of SwinV2-MoE, a MoE-based model built upon Swin Transformer V2 architecture. They reported superior accuracy in downstream computer vision tasks. A similar work (Puigcerver et al. 2023) introduced a soft MoE mechanism, demonstrating improved performance and training speed compared to classical MoE on billion-scale datasets.
Sparsely Activated MoE
@@
Based on conditional computations, MoE aims at activating specific parts of the model depending on the input. The core idea is to assign experts to different regions of the input space, thereby increasing the model capacity by augmenting the number of parameters without incurring significant computational overhead (Jacobs et al. 1991).
@@
According to Shazeer et al. (2017), an MoE includes a router $G$ and a set of experts $E$. The router learns a sparse assignment between the input and the experts, while the experts process the inputs like standard neural network modules do. Let $x$ be the input, and let $(E_{i})_{i\in[1,N]}$ be the $N$ experts of the MoE layer. Following  (Shazeer et al. 2017), the output of this MoE layer is given by:
$\sum_{i\in\text{Top}_{k}(x)}G(x)_{i}\cdot E_{i}(x),$
@@
where $G$ is a conv1x1 gate employing a softmax function, and $G(x)_{i}$ is its $i^{th}$ output. To promote sparsity in the MoE layer, the number of participating experts is restricted to $k<N$ for each input: $\text{Top}_{k}(x)$ contains the indices of the $k$ highest $G(x)_{i}$. Note that involving $k>1$ experts increases the computational cost compared to the sparse ($k=1$) counterpart. We tested several gate designs, but none of them clearly outperformed the simple conv1x1 gate (Table 10 in Appendix).
@@
However, enforcing this MoE design is insufficient to yield an effective MoE. To exploit compute capacity optimally, we encourage uniform usage of each expert by using a load balancing auxiliary loss proposed in (Shazeer et al. 2017), which promotes balanced expert assignments and equal importance of experts (For more details, see Section 4 of Shazeer et al. (2017)). We also use Batch Prioritized Routing (BPR) introduced by Riquelme et al. (2021).
Vision Transformer
@@
The Vision Transformer (ViT) (Dosovitskiy et al. 2020) represents a paradigm shift in computer vision. It moves away from traditional convolution-based architectures in favour of the Transformer’s attention mechanism, initially pioneered for natural language processing tasks (Vaswani et al. 2017). In the ViT model, an image is segmented into fixed-size and non-overlapping patches. These patches are linearly embedded into vectors and then passed through a series of Transformer blocks. Each Transformer block is predominantly made up of two main components: a multi-head self-attention mechanism and a multi-layer perceptron (MLP). Within the context of ViT, the MLP ratio (denoted mlp_ratio in the following) refers to the ratio of the hidden dimensions of the MLP to the embedding dimensions. This ratio is pivotal, as it determines the capacity and computational cost of the MLP component. We replace some predefined MLP block by a MoE, as depicted in Figure 1(a). Each MoE operates at the patch level: the experts are spatial distributed.
ConvNext
@@
ConvNext (Liu et al. 2022) has modernized the classical convolutional neural networks (CNN) design by incorporating insights from Vision Transformers (ViT). The central component of the ConvNext architecture is a block containing a $7\times 7$ depth-wise convolution followed by a $1\times 1$ convolutional layer (Figure 1(b)-left). This block is replicated across four stages, each with decreasing resolution. In this work, we replace some predefined ConvNext block, namely the multilayer perceptron, with a MoE block, and introduce a skip connection, as illustrated in Figure 1(b)-right. Consequently, each expert becomes a fully convolutional block operating at the feature map level, resulting in spatially distributed experts. Hence, this design does not only increase the network capacity, but it also allows each expert to specialize in specific spatial locations, and engages multiple experts per image.
Experiments on ImageNet
@@
This section presents the results obtained by the architectures described above (that incorporate MoE into the ConvNext and ViT architectures), trained on ImageNet-1k or pre-trained on ImageNet-21k datasets. It includes the experimental setup, the results on the ImageNet-1k validation set, and the impact of various MoE configurations.
Experimental setup
@@
All the results presented in this section rely on MoE models trained on the ImageNet datasets. Even though the context is different because of the changes of network structures presented in the previous sections, we set the training hyperparameters similar to (Touvron, Cord, and Jégou 2022) for ViT and (Liu et al. 2022) for ConvNext (a grid search looking for better settings did not bring significant improvement).
@@
Furthermore, when working with the ImageNet-1k dataset, we use a strong data-augmentation pipeline, including Mixup (Zhang et al. 2018), Cutmix (Yun et al. 2019), RandAugment (Cubuk et al. 2020), and Random Erasing (Zhong et al. 2020), over 300 epochs. Likewise, we utilize drop path, weight decay, and expert-specific weight decay as regularization strategies. However, for the pre-training phase on ImageNet-21k, we exclude Mixup and Random Erasing from the data augmentation pipeline because they did not improve results (as also reported in (Touvron, Cord, and Jégou 2022; Tu et al. 2022)). We pre-train the model for 90 epochs, adhering to the original ConvNext approach (Liu et al. 2022) and, consistently with Liu et al. (2022) and Touvron, Cord, and Jégou (2022), the final results of ImageNet-1k are obtained through fine-tuning for 30 epochs for ConvNext and 50 for ViT. Comprehensive details of all the hyperparameters are provided in Table 12 in the appendix.
Base results on ImageNet
@@
(×$\times 10^{6}$) #Paramsact (im/s) Accuracy
ConvNeXt-T (Liu et al. 2022) 28.6 28.6 4.5G 814 82.1
@@
Table 1: Accuracy for different ImageNet-1K trained models and the MoE strategies “Every 2” and “Last 2” (see text). “Top $k$” correspond to the number of experts involved. Throughput is measured on V100 GPUs, following (Touvron et al. 2021). For ImageNet-1k, non-isotropic ConvNext models feature an mlp_ratio of 2, which contributes to their improved flops-per-sample efficiency compared to their dense counterparts.
@@
(×$\times 10^{6}$) #Paramsact Accuracy Accuracy Accuracy
ConvNeXt-T (Liu et al. 2022) 28.6 28.6 4.5G 82.9 81.8 44.3
@@
Table 1 presents the results obtained on ImageNet-1k validation set by a model that has been entirely trained on ImageNet-1k, for isotropic architecture (e.g., ViT, ConvNext iso.) and a hierarchical architecture, namely ConvNext. We see some significant improvements in accuracy compared to the “no-MoE” results for small model size, especially for anisotropic models.
Table 2 presents the results of models that are pre-trained on ImageNet-21k, and tested on the same ImageNet-1k validation set than above. Here, MoE does bring some improvement for moderate numbers of activations per sample. However, this improvement decreases for large numbers of activations per sample. Figure 8 illustrates more examples.
Sensitivity with respect to design choices
This section presents different sensitivity studies that investigate the specifics of the design choices of the MoE layers. We explore the impact of the position of these layers inside the architecture, the effect of varying the number of experts, and finally, we assess the necessary design changes when transitioning from ImageNet-1k to ImageNet-21k dataset.
Impact of the position of MoE layers

### v2

# Mixture of Experts for Image Classification: What’s the Sweet Spot?

@@ Abstract
Mixture-of-Experts (MoE) models have shown promising potential for parameter-efficient scaling across domains. However, their application to image classification remains limited, often requiring billion-scale datasets to be competitive. In this work, we explore the integration of MoE layers into image classification architectures using open datasets. We conduct a systematic analysis across different MoE configurations and model scales. We find that moderate parameter activation per sample provides the best trade-off between performance and efficiency. However, as the number of activated parameters increases, the benefits of MoE diminish. Our analysis yields several practical insights for vision MoE design. First, MoE layers most effectively strengthen tiny and mid-sized models, while gains taper off for large-capacity networks and do not redefine state-of-the-art ImageNet performance. Second, a Last-2 placement heuristic offers the most robust cross-architecture choice, with Every-2 slightly better for Vision Transform (ViT), and both remaining effective as data and model scale increase. Third, larger datasets (e.g., ImageNet-21k) allow more experts, up to 16, for ConvNeXt to be utilized effectively without changing placement, as increased data reduces overfitting and promotes broader expert specialization. Finally, a simple linear router performs best, suggesting that additional routing complexity yields no consistent benefit.
1 Introduction
@@
Recent advances in machine learning, particularly in the domains of natural language processing (NLP)(Vaswani et al., 2017; Kenton & Toutanova, 2019) and computer vision (Dosovitskiy et al., 2020), have been primarily driven by scaling up model size, computational budgets, and training data. Although these large-scale models demonstrate impressive performance, they are often expensive to train and consume considerable energy resources (Strubell et al., 2019). As a result, the research community has become increasingly interested in exploring more efficient training and serving paradigms. One such promising solution is the use of sparse expert models, with Mixture-of-Experts (MoE) (Shazeer et al., 2017; Lepikhin et al., 2020) emerging as a popular variant.
@@
MoE models introduce sparsity by partitioning the set of parameters into multiple parallel sub-models, called “experts.” During training and inference, the gating part of the models routes input examples to specific expert(s), ensuring that each example only interacts with a subset of the network parameters. As the computational cost is partially correlated with the number of parameters activated for a given sample rather than the total number of parameters, this approach facilitates the scaling-up of the model, while keeping computational costs under control, making it an attractive option for a wide range of applications.
@@
Despite their success in various domains (Hihn & Braun, 2021; Costa-jussà et al., 2022; Zoph et al., 2022; Fedus et al., 2022), the application of MoE models in image classification and computer vision, in general, remains limited, and it often requires very large datasets (Riquelme et al., 2021; Mustafa et al., 2022; Komatsuzaki et al., 2022) to be competitive against state-of-the-art approaches. In this work, we focus on leveraging the potential of MoE models for image classification on ImageNet-1k and ImageNet-21k (Russakovsky et al., 2015). We study the efficiency of integrating MoE within two renowned architectures, ConvNext (Liu et al., 2022) and Vision Transformer (ViT) (Touvron et al., 2022). We conduct a series of experiments considering various architecture configurations. Likewise, we investigate the impact of various components, including the number of experts and their sizes, the gate design, and the layer positions, among others. Our experimental findings indicate that optimal design is contingent on the specific network architecture, and that consistently situating the MoE layer within the final two even blocks invariably yields substantial improvements for moderate model size. Nevertheless, when scaling up the approach to large models and datasets close to the state-of-the-art, the benefits of using Mixture-of-Experts for image classification gradually vanish.
Contributions.
In this work, we provide a systematic study of applying Mixture-of-Experts (MoE) models to image classification tasks, focusing on the ImageNet-1k and ImageNet-21kt benchmarks. Training follows supervised image classification in two setups, either on ImageNet-1K or ImageNet-22K. Unlike prior efforts that emphasize large-scale deployments with vast compute resources, we explore MoE integration within mid-sized ConvNeXt and Vision Transformer (ViT) architectures, identifying effective design principles for efficient training and inference. Our analysis yields the following practical guidance for vision MoE design:
  • When MoE helps (scale): MoE strengthens tiny and mid-sized models, but gains diminish for large-capacity networks, and do not redefine state-of-the-art ImageNet performance.
  • Placement heuristic: A Last-2 configuration is the most robust cross-architecture choice, with Every-2 slightly better for ViT. These placements remain effective when scaling data or model size.
  • Scaling with data: With larger datasets (e.g., ImageNet-21k), more experts can be used effectively, up to 16 for ConvNeXt, without changing placement. Increased data reduces overfitting and enables broader expert utilization.
  • Routing choice: A simple linear router works the best, indicating that added routing complexity brings no consistent benefit.
2 Related Work
@@
The Mixture-of-Experts (MoE) model, introduced by (Jacobs et al., 1991), partitions complex tasks into hopefully simpler sub-tasks handled by expert models, whose predictions are combined to produce the final output. This framework has been successfully integrated into various neural network architectures, particularly transformers for NLP tasks (Shazeer et al., 2017). Given this success in NLP, interest has increased in applying MoE to the diverse and complex computer vision tasks.
@@
The transformative potential of MoE architectures is underscored by their successful integration across various domains, with the Vision Transformer (ViT) being a prime example. Riquelme et al. (2021) introduced V-MoE, an MoE-augmented ViT model, for image classification tasks on a massive dataset containing hundreds of millions of examples. They showed that V-MoE not only matches the performance of prior state-of-the-art architectures but also requires half the computational resources during inference. Lou et al. (2021) presented a sparse MoE MLP model based on the MLP-Mixer architecture (Tolstikhin et al., 2021). While the MLP-Mixer architecture does not display performance accuracy, the MoE-enhanced version surpassed its dense counterpart in experiments on ImageNet and CIFAR. Hwang et al. (2023a) demonstrated the effectiveness of SwinV2-MoE, a MoE-based model built upon Swin Transformer V2 architecture. They reported superior accuracy in downstream computer vision tasks. Similarly, Puigcerver et al. (2024) introduced a soft MoE mechanism, demonstrating improved performance and training speed compared to classical MoE on billion-scale datasets. More recently, Han et al. (2024) proposed ViMoE, which investigates expert placement strategies by fine-tuning DINOv2 models. In contrast, our work focuses on training MoE architectures from scratch, enabling a systematic study of their behavior across both ConvNeXt and ViT backbones. This distinction allows us to provide complementary insights into how MoE mechanisms contribute to compute efficiency and performance in vision models.
3 Sparsely Activated MoE
@@
Based on conditional computations, MoE aims at activating specific parts of the model depending on the input. The core idea is to assign experts to different regions of the input space, thereby increasing the model capacity by augmenting the number of parameters without incurring significant computational overhead (Jacobs et al., 1991).
@@
According to Shazeer et al. (2017), an MoE includes a router $G$ and a set of experts $E$. The router learns a sparse assignment between the input and the experts, while the experts process the inputs like standard neural network modules do. Let $x$ be the input, and let $(E_{i})_{i\in[1,N]}$ be the $N$ experts of the MoE layer. Following (Shazeer et al., 2017), the output of this MoE layer is given by:
$\sum_{i\in\text{Top}_{k}(x)}G(x)_{i}\cdot E_{i}(x)$
@@
where, $G$ is a conv1x1 gate employing a softmax function, and $G(x)_{i}$ is its $i^{th}$ output. To promote sparsity in the MoE layer, the number of participating experts is restricted to $k<N$ for each input: $\text{Top}_{k}(x)$ contains the indices of the $k$ highest $G(x)_{i}$. Note that involving $k>1$ experts increases the computational cost compared to the sparse ($k=1$) counterpart. We tested several gate designs, but none of them clearly outperformed the simple conv1x1 gate (Sec.˜4.3.5).
@@
However, enforcing this MoE design is insufficient to yield an effective MoE. To exploit compute capacity optimally, we encourage uniform usage of each expert by using a load balancing auxiliary loss proposed in (Shazeer et al., 2017), which promotes balanced expert assignments and equal importance of experts (for more details, see Section 4 of Shazeer et al. (2017)). We also use Batch Prioritized Routing (BPR) introduced by (Riquelme et al., 2021).
3.1 Vision Transformer
@@
The Vision Transformer (ViT) (Dosovitskiy et al., 2020) represents a paradigm shift in computer vision. It moves away from traditional convolution-based architectures in favor of the Transformer’s attention mechanism, initially pioneered for natural language processing tasks (Vaswani et al., 2017). In the ViT model, an image is segmented into fixed-size and non-overlapping patches. These patches are linearly embedded into vectors and then passed through a series of Transformer blocks. Each Transformer block is predominantly made up of two main components: a multi-head self-attention mechanism and a multi-layer perceptron (MLP). Within the context of ViT, the MLP ratio (denoted mlp_ratio in the following) refers to the ratio of the hidden dimensions of the MLP to the embedding dimensions. This ratio is pivotal, as it determines the capacity and computational cost of the MLP component. We replace some predefined MLP block by a MoE, as depicted in Fig.˜1(a). Each MoE operates at the patch level: the experts are spatially distributed. We consider (a) a standard ViT architecture (ViT-B (Base)) and (b) a variant with 8 experts (ViT-B-8).
3.2 ConvNeXt
@@
ConvNeXt (Liu et al., 2022) has modernized the classical convolutional neural networks (CNN) design by incorporating insights from Vision Transformers (ViT). The central component of the ConvNeXt architecture is a block containing a $7\times 7$ depth-wise convolution followed by a $1\times 1$ convolutional layer (Fig.˜1(b)-left). This block is replicated across four stages, each with decreasing resolution. In this work, we replace some predefined ConvNeXt block, namely the multilayer perceptron, with a MoE block, and introduce a skip connection, as illustrated in Fig.˜1(b)-right. Consequently, each expert becomes a fully convolutional block operating at the feature map level, resulting in spatially distributed experts. Hence, this design not only increases the network capacity but also allows each expert to specialize in specific spatial locations and engages multiple experts per image. We examine three variants: (a) ConvNeXt-T (Tiny), ConvNeXt-S (Small), and ConvNeXt-B (Base). These architectures primarily differ in depth and width; and (b) ConvNeXt-S-4 and ConvNeXt-B-4, which are variants with 4 experts.
4 Experiments on ImageNet
@@
This section presents the results obtained by the architectures described above (that incorporate MoE into ConvNeXt and ViT architectures), trained on ImageNet-1k or pre-trained on ImageNet-21k datasets. It includes the experimental setup, the results on the ImageNet-1k validation set, and the impact of various MoE configurations.
4.1 Experimental Setup
@@
All the results presented in this section rely on MoE models trained on the ImageNet datasets. Although the context is different due to the changes in network structures presented in Secs.˜3.1 and 3.2, we set the training hyperparameters similar to (Touvron et al., 2022) for ViT and (Liu et al., 2022) for ConvNeXt (a grid search looking for better settings did not bring significant improvement).
@@
Furthermore, when working with the ImageNet-1k dataset, we use a strong data-augmentation pipeline, including Mixup (Zhang et al., 2018), Cutmix (Yun et al., 2019), RandAugment (Cubuk et al., 2020), and Random Erasing (Zhong et al., 2020), over 300 epochs. Likewise, we utilize drop path, weight decay, and expert-specific weight decay as regularization strategies. However, for the pretraining phase on ImageNet-21k, we exclude Mixup and Random Erasing from the data augmentation pipeline because they did not improve the results, as also reported in (Touvron et al., 2022; Tu et al., 2022). We pre-train the model for 90 epochs, adhering to the original ConvNeXt approach (Liu et al., 2022) and, consistently with (Liu et al., 2022) and (Touvron et al., 2022), the final results of ImageNet-1k are obtained through fine-tuning for 30 epochs for ConvNeXt and 50 for ViT. Comprehensive details of all the hyperparameters are provided in Tab.˜10 in App.˜A.
4.2 Base results on ImageNet
Table 1: Accuracy for different ImageNet-1K trained models and the MoE strategies “Every 2” and “Last 2” (see text). “Top $k$” corresponds to the number of experts involved. Throughput is measured on V100 GPUs, following (Touvron et al., 2021). For ImageNet-1k, non-isotropic ConvNeXt models feature an mlp_ratio of 2, which contributes to their improved flops-per-sample efficiency compared to their dense counterparts.
@@ 4.2 Base results on ImageNet
($\times 10^{6}$) #Paramsact (im/s) Accuracy
ConvNeXt-T (Liu et al., 2022) 28.6 28.6 4.5G 814 82.1
@@
($\times 10^{6}$) #Paramsact Accuracy Accuracy Accuracy
ConvNeXt-T (Liu et al., 2022) 28.6 28.6 4.5G 82.9 81.8 44.3
@@
Tab.˜12 presents the results obtained on ImageNet-1k validation set by a model that has been entirely trained on ImageNet-1k, for isotropic architecture (e.g., ViT, ConvNeXt iso.) and a hierarchical architecture, namely ConvNeXt. We see some significant improvements in accuracy compared to the “no-MoE” results for small model sizes, especially for anisotropic models.
Tab.˜2 presents the results of models that are pre-trained on ImageNet-21k, and tested on the same ImageNet-1k validation set than above. Here, MoE does bring some improvement for moderate numbers of activations per sample. However, this improvement decreases for large numbers of activations per sample, as depicted in Fig.˜2.
Figure 2: Left: Pareto-front for ImageNet21k, x-axis = number of activations per sample. ViT models have been presented in MoE versions only after additional pretraining, and are therefore not presented. MoE seems to be Pareto optimal for a number of activations per sample below 90M. Right: Pareto-front for ViT models on JFT-300M. Overall, MoE is never validated for a number of activations per sample above $\approx$ 100M.
MoE Strengthens Moderate Model Size; Not the Frontier MoE integration provides clear gains at a moderate scale, with tiny and small models benefiting most. As data scale grows, these relative improvements strengthen. Yet as model capacity rises, these benefits diminish, showing that MoE enhances mid-sized models but does not redefine SoTA performance.
4.3 Sensitivity with Respect to Design Choices
This section presents different sensitivity studies that investigate the specifics of the design choices of the MoE layers. We explore the impact of the position of these layers inside the architecture (Sec.˜4.3.1), the effect of varying the number of experts (Sec.˜4.3.2), and finally, we assess the necessary design changes when transitioning from the ImageNet-1k to ImageNet-21k dataset (Sec.˜4.3.4) and the impact of the routing architecture choices (Sec.˜4.3.5).Ideally, sensitivity analyses would explore the joint effects of multiple factors (such as the number, placement, and size of experts), but exhaustive multi-axis exploration quickly becomes computationally intractable. Nevertheless, we partially address such interactions by analyzing the combined impact of expert size and expert count in LABEL:sec:abla:size_number, highlighting how these factors jointly influence efficiency and accuracy.
4.3.1 Impact of the Position of MoE Layers